How do I solve the following by factoring?:
X^2 + 7x +10 = 0

Answer by nerdybill(7384)

X^2 + 7x +10 = 0
.
(x+2)(x+5) = 0
.
Now, if either of the two terms on the left is zero, the entire side will be zero.
.
Two solutions, can be found:
(x+2) = 0
x = -2
.
(x+5) = 0
x = -5
.
Therefore,
x = {-5, -2}